# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ 4. Characterization of Rate Function
$\Big{|}\int_{0}^{T}\left(\left\langle R_{2}\theta_{n},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta_{n}\right\rangle-\left\langle R_{1}\theta_{n},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta_{n}\right\rangle\right)\mathrm{d}t$
$-\int_{0}^{T}\left(\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle-\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\right)\mathrm{d}t$ $\Big{|}\leqslant I^{n}_{1j}+I^{n}_{2j}+I^{n}_{3j}+I_{4j},$
where
$I^{n}_{1j}$ $:=\left|\int_{0}^{T}\langle\nabla\cdot\left(\theta_{n}R^{\perp}\left(S_{j}\theta_{n}\right)\right),\varphi\rangle\,\mathrm{d}t-\int_{0}^{T}\langle\nabla\cdot\left(\theta R^{\perp}\left(S_{j}\theta\right)\right),\varphi\rangle\,\mathrm{d}t\right|,$
$I^{n}_{2j}$ $:=\left|\int_{0}^{T}\langle\nabla\cdot\left(S_{j}\theta_{n}R^{\perp}\left(H_{j}\theta_{n}\right)\right),\varphi\rangle\,\mathrm{d}t-\int_{0}^{T}\langle\nabla\cdot\left(S_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),\varphi\rangle\,\mathrm{d}t\right|,$
$I^{n}_{3j}$ $:=\left|\int_{0}^{T}\langle\nabla\cdot\left(H_{j}\theta_{n}R^{\perp}\left(H_{j}\theta_{n}\right)\right),\varphi\rangle\,\mathrm{d}t\right|,$
and
$I_{4j}:=\left|\int_{0}^{T}\langle\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),\varphi\rangle\,\mathrm{d}t\right|.$
The convergence of $\theta_{n}$ in $\mathbb{X}_{\alpha,\beta}$ implies that the $L^{2}\left([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2})\right)$-norm of $\{\theta_{n}\}$ is bounded:
$\int_{0}^{T}\left\|\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t\leqslant C.$ (4.15)
Using (2.4) for $r=2\beta-2\alpha$, (4.15) and the fact that $\left\|H_{j}\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}\leqslant\left\|\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})},$ we have
@@
$\lesssim C(\varphi)2^{-j(4\alpha+1-4\beta)}\int_{0}^{T}\left\|\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t$
$\lesssim C(\varphi)2^{-j(4\alpha+1-4\beta)}.$
@@
According to (4.16) and (4.17), for any $\eta>0,$ we can choose a sufficiently large positive integer $j_{0},$ such that $I^{n}_{3j_{0}}+I_{4j_{0}}\leqslant\eta$ for all $n\geqslant 1.$ By [Mar08a, Lemma 9.3], $S_{j_{0}}\theta_{n}\rightarrow S_{j_{0}}\theta$ and $R^{\perp}(S_{j_{0}}\theta_{n})\rightarrow R^{\perp}(S_{j_{0}}\theta)$ strongly in $L^{2}([0,T];H^{1/2}(\mathbb{T}^{2}))$, which implies
@@
5. Upper Bound for Large Derivations
@@
For every $\varepsilon,\delta>0$ and $m\in\mathbb{N}_{+}$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A, and let $\theta_{\varepsilon,m}$ be a solution of (3.7) with initial data $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$.
(i) Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta}$ be the laws of $\theta_{\varepsilon,\delta}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.7) holds for $(\varepsilon,\delta)$, then for any closed set $F\subset\mathbb{X}_{\alpha,\beta}$,
@@
(ii) Let $\mu_{\varepsilon}=\mu_{\varepsilon,m}$ be the laws of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.8) holds for $(\varepsilon,m)$, then for any closed set $F\subset\mathbb{X}_{\alpha,\beta}$, the upper bound (5.1) holds as well.
@@
By Assumption A, Proposition B.1 and Varadhan’s integration lemma (see, for example, [DZ10, Theorem 4.3.1]),
@@
Since $\theta_{\varepsilon,\delta}$ is a solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$, the stochastic process $\mathcal{M}^{\psi,\varphi}(t,\theta_{\varepsilon,\delta})$ is a martingale. Due to the contraction property of $Q_{\delta}$ on $H^{2\beta}(\mathbb{T}^{2})$, the quadratic variation of $\mathcal{M}^{\psi,\varphi}(\cdot,\theta_{\varepsilon,\delta})$ satisfies
$\langle\langle\mathcal{M}^{\psi,\varphi}(\cdot,\theta_{\varepsilon,\delta})\rangle\rangle_{t}=\varepsilon\int_{0}^{t}\|\sqrt{Q_{\delta}}\Lambda^{2\beta}\varphi\|_{H}^{2}\,\mathrm{d}s\leqslant\varepsilon\int_{0}^{t}\|\varphi\|_{H^{2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
It follows that $Q^{\psi,\varphi}(t,\theta_{\varepsilon,\delta})$ is a supermartingale. For every compact set $F\subset\mathbb{X}_{\alpha,\beta}$, we have
@@
Since $F$ is compact in $\mathbb{X}_{\alpha,\beta}$, by [KL99, Appendix 2, Lemma 3.2], one can swap the above infimum and supremum. Finally, Proposition 4.2 implies that
@@
6. Lower Bound for Large Derivations
In this section, we will prove a restricted large deviations lower bound using the entropy method as described in [Mar10] (see Lemma 2.5). Due to the lack of well-posedness of the skeleton equation (1.1), it is challenging to match the upper and lower bounds in $\mathbb{X}_{\alpha,\beta}$. Instead, we will match the upper and lower bounds within the following weak-strong uniqueness class.
Definition 6.1 (Weak-Strong Uniqueness Class).
We say $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ is a weak-strong uniqueness class of (1.1) for every $g\in L^{2}([0,T];H)$ if, for arbitrary two weak solutions $\theta_{1},\theta_{2}$ of the skeleton equation (1.1) in the sense of Definition 4.1 with the same initial data $\theta_{1}(0)=\theta_{2}(0)$ and the same control $g$, we have $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\beta}$ as long as $\theta_{1}\in\mathcal{C}_{0}$ and $\theta_{2}$ satisfies the $H^{\alpha-2\beta}$-energy inequality: for every $t\in[0,T],$
Definition 6.2 ($\mathcal{I}$-closure).
Suppose that $\mathcal{A}$ is a subset of $\mathbb{X}_{\alpha,\beta}$ and $\mathcal{I}:\mathbb{X}_{\alpha,\beta}\rightarrow[0,+\infty]$ is lower semi-continuous. We define the $\mathcal{I}$-closure of $\mathcal{A}$ as
$\overline{\mathcal{A}}^{\mathcal{I}}:=\big{\{}\theta\in\mathbb{X}_{\alpha,\beta}:\text{there exists a sequence }\{\theta^{(n)}\}\subset\mathcal{A},\,\text{such that }\theta^{(n)}\rightarrow\theta\text{ in }\mathbb{X}_{\alpha,\beta},\,\mathcal{I}\big{(}\theta^{(n)}\big{)}\rightarrow\mathcal{I}(\theta)\big{\}}.$
@@ 6. Lower Bound for Large Derivations
For every $\varepsilon,\delta>0$ and $m\in\mathbb{N}_{+}$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A, and let $\theta_{\varepsilon,m}$ be a solution of (3.7) with initial data $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$. Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) and let $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$.
(i) Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta}$ be the law of $\theta_{\varepsilon,\delta}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.7) holds for $(\varepsilon,\delta)$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$,
@@
(ii) Let $\mu_{\varepsilon}=\mu_{\varepsilon,m}$ be the law of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.8) holds for $(\varepsilon,m)$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$, the restricted lower bound (6.2) holds as well.
@@
For every $\theta\in\mathcal{C}_{0}$ with $\mathcal{I}(\theta)<\infty$, let $g\in L^{2}([0,T];H)$ be the control such that $\theta$ solves (1.1) weakly. For every $\varepsilon>0$ and $m\in\mathbb{N}_{+}$ satisfying the scaling regime (2.8), let $\theta_{\varepsilon,m}$ and $\mu_{\varepsilon,m}$ be as in Proposition 6.3. Then there exists a family of probability measure $\{\pi_{\varepsilon,m,\theta}\}$ on $\mathbb{X}_{\alpha,\beta}$ that satisfies the entropy inequality
$\limsup_{\varepsilon\rightarrow 0}\operatorname{Ent}\left(\pi_{\varepsilon,m,\theta}\mid\mu_{\varepsilon,m}\right)\leqslant\mathcal{I}(\theta).$ (6.3)
@@
Let $\left(\Omega,\mathcal{F},\{\mathcal{F}(t)\}_{t\in[0,T]},\overline{\mathbb{P}}\right)$ be a stochastic basis and $\overline{W}$ be a cylindrical Wiener process on $H^{\alpha-2\beta}(\mathbb{T}^{2})$ such that $\bar{\theta}_{\varepsilon,m}$ is a solution of (3.7) with initial data $\bar{\theta}_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$ and $\bar{\theta}_{\varepsilon,m}(0)$ is independent of $\overline{W}$. For any two measures $\mu,\nu$ on $\Omega$, we denote by $\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$ the Radon-Nikodym derivative of $\mu$ with respect to $\nu.$ Now we define the $H$-valued random variable $Y_{0}^{\varepsilon,m}$ by
$Y_{0}^{\varepsilon,m}:=\frac{\mathrm{d}\mathcal{G}(P_{m}\theta(0),\varepsilon P_{m}/2)}{\mathrm{d}\mathcal{G}(0,\varepsilon P_{m}/2)}\left(\bar{\theta}_{\varepsilon,m}(0)\right),$
and define the stochastic process $Z^{\varepsilon,m}$ by
$Z^{\varepsilon,m}(t):=\exp\left(-\frac{1}{\sqrt{\varepsilon}}\int_{0}^{t}\left\langle P_{m}g(s),\,\mathrm{d}\overline{W}(s)\right\rangle-\frac{1}{2\varepsilon}\int_{0}^{t}\left\|P_{m}g(s)\right\|_{H}^{2}\,\mathrm{d}s\right)Y_{0}^{\varepsilon,m},\quad t\in[0,T].$ (6.4)
Let $\mathbb{P}$ be the probability measure on $\left(\Omega,\mathcal{F}\right)$ such that $Z^{\varepsilon,m}(T)=\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\overline{\mathbb{P}}}.$ By Cameron-Martin theorem, $W(t):=\overline{W}(t)+\varepsilon^{-1/2}\int_{0}^{t}P_{m}g(s)\,\mathrm{d}s$ is a cylindrical Wiener process on $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ and $\bar{\theta}_{\varepsilon,m}$ is a solution of
$\mathrm{d}\bar{\theta}_{\varepsilon,m}(t)=-\Lambda^{2\alpha}\bar{\theta}_{\varepsilon,m}(t)\,\mathrm{d}t-P_{m}\left(u_{\bar{\theta}_{\varepsilon,m}}(t)\cdot\nabla\bar{\theta}_{\varepsilon,m}(t)\right)\,\mathrm{d}t+\sqrt{\varepsilon}\Lambda^{2\beta}\,\mathrm{d}W^{m}(t)+P_{m}\Lambda^{2\beta}g\,\mathrm{d}t$ (6.5)
with the initial data $\bar{\theta}_{\varepsilon,m}(0)\sim\mathcal{G}(P_{m}\theta(0),\varepsilon P_{m}/2).$
Let $\pi_{\varepsilon,m,\theta}$ be the law of $\bar{\theta}_{\varepsilon,m}$ on $\left(\Omega,\mathcal{F},\mathbb{P}\right)$. According to (6.4), the definition of relative entropy, the expression of $\mathcal{I}_{0}$, and the proof for the lower bound of Proposition B.1,
$\varepsilon\operatorname{Ent}\left(\pi_{\varepsilon,m,\theta}|\mu_{\varepsilon,m}\right)\leqslant\varepsilon\operatorname{Ent}(\mathbb{P}|\overline{\mathbb{P}})$ $=\varepsilon\mathbb{E}\left[\log Z^{\varepsilon,m}(T)\right]$
$=\frac{1}{2}\int_{0}^{T}\|P_{m}g(s)\|_{H}^{2}\,\mathrm{d}s+\varepsilon\mathbb{E}\left[\log Y_{0}^{\varepsilon,m}\right]$
@@
For every $\varepsilon>0$ and $m\in\mathbb{N}_{+}$ satisfying the scaling regime (2.8), let $\theta,g,\pi^{\varepsilon,m,\theta}$ and $\mu_{\varepsilon,m}$ be as in Lemma 6.4. Then $\pi^{\varepsilon,m,\theta}$ converges to $\delta_{\theta}$ in law on $\mathbb{X}_{\alpha,\beta}$ as $\varepsilon\rightarrow 0$.
@@
The proof will be divided into three steps. First, we show that $\{\pi^{\varepsilon,m,\theta}\}$ is tight in $\mathbb{X}_{\alpha,\beta}$ and that any limit point $\bar{\theta}$ is a weak solution of (1.1) in the sense of Definition 4.1. Next, we prove the $H^{\alpha-2\beta}$-energy inequality (6.1) for $\bar{\theta}$. Finally, we conclude that $\bar{\theta}=\theta$ in $\mathbb{X}_{\alpha,\beta}$ by using the weak-strong uniqueness.
Step 1. $\bar{\theta}$ is a weak solution of the skeleton equation.

### v3

@@ 4. Characterization of Rate Function
$\quad\Big|\int_{0}^{T}\left(\left\langle R_{2}\theta_{n},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta_{n}\right\rangle-\left\langle R_{1}\theta_{n},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta_{n}\right\rangle\right)\mathrm{d}t$
$\quad\quad-\int_{0}^{T}\left(\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle-\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\right)\mathrm{d}t\Big|$
$\leqslant\left|\int_{0}^{T}\langle\nabla\cdot\left(\theta_{n}R^{\perp}\left(S_{j}\theta_{n}\right)\right),\varphi\rangle\,\mathrm{d}t-\int_{0}^{T}\langle\nabla\cdot\left(\theta R^{\perp}\left(S_{j}\theta\right)\right),\varphi\rangle\,\mathrm{d}t\right|$
$+\left|\int_{0}^{T}\langle\nabla\cdot\left(S_{j}\theta_{n}R^{\perp}\left(H_{j}\theta_{n}\right)\right),\varphi\rangle\,\mathrm{d}t-\int_{0}^{T}\langle\nabla\cdot\left(S_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),\varphi\rangle\,\mathrm{d}t\right|$
$+\left|\int_{0}^{T}\langle\nabla\cdot\left(H_{j}\theta_{n}R^{\perp}\left(H_{j}\theta_{n}\right)\right),\varphi\rangle\,\mathrm{d}t\right|+\left|\int_{0}^{T}\langle\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),\varphi\rangle\,\mathrm{d}t\right|$
$=:I^{n}_{1j}+I^{n}_{2j}+I^{n}_{3j}+I_{4j}.$
The convergence of $\theta_{n}$ in $\mathbb{X}_{\alpha,\beta}$ implies the boundedness of $L^{2}\left([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2})\right)$-norm of $\{\theta_{n}\}$. Combining this with (2.4) for $r=2\beta-2\alpha$ and the fact that $\left\|H_{j}\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}\leqslant\left\|\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})},$ we have
@@
$\lesssim C(\varphi)2^{-j(4\alpha+1-4\beta)}\int_{0}^{T}\left\|\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t\lesssim C(\varphi)2^{-j(4\alpha+1-4\beta)}.$
@@
According to (4.15) and (4.16), for any $\eta>0,$ we can choose a sufficiently large positive integer $j_{0},$ such that $I^{n}_{3j_{0}}+I_{4j_{0}}\leqslant\eta$ for all $n\geqslant 1.$ By [Mar08_existence, Lemma 9.3], $S_{j_{0}}\theta_{n}\rightarrow S_{j_{0}}\theta$ and $R^{\perp}(S_{j_{0}}\theta_{n})\rightarrow R^{\perp}(S_{j_{0}}\theta)$ strongly in $L^{2}([0,T];H^{1/2}(\mathbb{T}^{2}))$, which implies
@@
5. Upper Bound for Large Deviations
@@
For every $\varepsilon,\delta(\varepsilon)>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$, let $\theta_{\varepsilon,\delta(\varepsilon)}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 1.2 with initial data $\theta_{\varepsilon,\delta(\varepsilon)}(0)$ satisfying Assumption A, and let $\theta_{\varepsilon,m(\varepsilon)}$ be the solution of (3.1) with initial data $\theta_{\varepsilon,m(\varepsilon)}(0)\sim\mathcal{G}(0,\varepsilon P_{m(\varepsilon)}/2)$.
(i) Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta(\varepsilon)}$ be the laws of $\theta_{\varepsilon,\delta(\varepsilon)}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (1.7) holds for $(\varepsilon,\delta(\varepsilon))$, then for any closed set $F\subset\mathbb{X}_{\alpha,\beta}$,
@@
(ii) Let $\mu_{\varepsilon}=\mu_{\varepsilon,m(\varepsilon)}$ be the laws of $\theta_{\varepsilon,m(\varepsilon)}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.5) holds for $(\varepsilon,m(\varepsilon))$, then for any closed set $F\subset\mathbb{X}_{\alpha,\beta}$, the upper bound (5.1) holds as well.
@@
By Assumption A, Proposition 2.4, and Varadhan’s integration lemma (see, for example, [DZ10_LDP, Theorem 4.3.1]),
@@
Since $\theta_{\varepsilon,\delta(\varepsilon)}$ is a solution of (1.3) in the sense of Definition 1.2 with initial data $\theta_{\varepsilon,\delta(\varepsilon)}(0)$, the stochastic process $\mathcal{M}^{\psi,\varphi}(t,\theta_{\varepsilon,\delta(\varepsilon)})$ is a martingale. Due to the contraction property of $Q_{\delta(\varepsilon)}$ on $H^{2\beta}(\mathbb{T}^{2})$, the quadratic variation of $\mathcal{M}^{\psi,\varphi}(\cdot,\theta_{\varepsilon,\delta(\varepsilon)})$ satisfies
$\langle\langle\mathcal{M}^{\psi,\varphi}(\cdot,\theta_{\varepsilon,\delta(\varepsilon)})\rangle\rangle_{t}=\varepsilon\int_{0}^{t}\|\sqrt{Q_{\delta(\varepsilon)}}\Lambda^{2\beta}\varphi\|_{H}^{2}\,\mathrm{d}s\leqslant\varepsilon\int_{0}^{t}\|\varphi\|_{H^{2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
It follows that $Q^{\psi,\varphi}(t,\theta_{\varepsilon,\delta(\varepsilon)})$ is a supermartingale. For every compact set $F\subset\mathbb{X}_{\alpha,\beta}$, we have
@@
Since $F$ is compact in $\mathbb{X}_{\alpha,\beta}$, by [KL99_IPS, Appendix 2, Lemma 3.2], one can swap the above infimum and supremum. Finally, Proposition 4.2 implies that
@@
6. Lower Bound for Large Deviations
In this section, we will prove a restricted large deviations lower bound using the entropy method as described in [Mariani10] (see Lemma 2.3). Due to the lack of well-posedness of the skeleton equation (1.1), it is challenging to match the upper and lower bounds in $\mathbb{X}_{\alpha,\beta}$. Instead, we will match the upper and lower bounds within the $\mathcal{I}$-closure of the weak-strong uniqueness class defined in Definition 1.1.
@@ 6. Lower Bound for Large Deviations
For every $\varepsilon,\delta(\varepsilon)>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$, let $\theta_{\varepsilon,\delta(\varepsilon)}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 1.2 with initial data $\theta_{\varepsilon,\delta(\varepsilon)}(0)$ satisfying Assumption A, and let $\theta_{\varepsilon,m(\varepsilon)}$ be the solution of (3.1) with initial data $\theta_{\varepsilon,m(\varepsilon)}(0)\sim\mathcal{G}(0,\varepsilon P_{m(\varepsilon)}/2)$. Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) and let $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$.
(i) Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta(\varepsilon)}$ be the law of $\theta_{\varepsilon,\delta(\varepsilon)}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (1.7) holds for $(\varepsilon,\delta(\varepsilon))$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$,
@@
(ii) Let $\mu_{\varepsilon}=\mu_{\varepsilon,m(\varepsilon)}$ be the law of $\theta_{\varepsilon,m(\varepsilon)}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.5) holds for $(\varepsilon,m(\varepsilon))$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$, the restricted lower bound (6.1) holds as well.
@@
For every $\theta\in\mathcal{C}_{0}$ with $\mathcal{I}(\theta)<\infty$, let $g\in L^{2}([0,T];H)$ be the control such that $\theta$ solves (1.1) weakly. For every $\varepsilon>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$ satisfying the scaling regime (2.5), let $\theta_{\varepsilon,m(\varepsilon)}$ and $\mu_{\varepsilon,m(\varepsilon)}$ be as in Proposition 6.1. Then there exists a family of probability measure $\{\pi_{\varepsilon,m(\varepsilon),\theta}\}$ on $\mathbb{X}_{\alpha,\beta}$ that satisfies the entropy inequality
$\limsup_{\varepsilon\rightarrow 0}\operatorname{Ent}\left(\pi_{\varepsilon,m(\varepsilon),\theta}\mid\mu_{\varepsilon,m(\varepsilon)}\right)\leqslant\mathcal{I}(\theta).$ (6.2)
@@
Let $\left(\Omega,\mathcal{F},\{\mathcal{F}(t)\}_{t\in[0,T]},\overline{\mathbb{P}}\right)$ be a stochastic basis and $\overline{W}$ be a cylindrical Wiener process on $H^{\alpha-2\beta}(\mathbb{T}^{2})$ such that $\bar{\theta}_{\varepsilon,m(\varepsilon)}$ is a solution of (3.1) with initial data $\bar{\theta}_{\varepsilon,m(\varepsilon)}(0)\sim\mathcal{G}(0,\varepsilon P_{m(\varepsilon)}/2)$ and $\bar{\theta}_{\varepsilon,m(\varepsilon)}(0)$ is independent of $\overline{W}$. For any two measures $\mu,\nu$ on $\Omega$, we denote by $\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$ the Radon-Nikodym derivative of $\mu$ with respect to $\nu.$ Now we define the $H$-valued random variable $Y_{0}^{\varepsilon,m(\varepsilon)}$ by
$Y_{0}^{\varepsilon,m(\varepsilon)}:=\frac{\mathrm{d}\mathcal{G}(P_{m(\varepsilon)}\theta(0),\varepsilon P_{m(\varepsilon)}/2)}{\mathrm{d}\mathcal{G}(0,\varepsilon P_{m(\varepsilon)}/2)}\left(\bar{\theta}_{\varepsilon,m(\varepsilon)}(0)\right),$
and define the stochastic process $Z^{\varepsilon,m(\varepsilon)}$ by
$Z^{\varepsilon,m(\varepsilon)}(t):=\exp\left(-\frac{1}{\sqrt{\varepsilon}}\int_{0}^{t}\left\langle P_{m(\varepsilon)}g(s),\,\mathrm{d}\overline{W}(s)\right\rangle-\frac{1}{2\varepsilon}\int_{0}^{t}\left\|P_{m(\varepsilon)}g(s)\right\|_{H}^{2}\,\mathrm{d}s\right)Y_{0}^{\varepsilon,m(\varepsilon)}.$ (6.3)
Let $\mathbb{P}$ be the probability measure on $\left(\Omega,\mathcal{F}\right)$ such that $Z^{\varepsilon,m(\varepsilon)}(T)=\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\overline{\mathbb{P}}}.$ By Cameron-Martin theorem, $W(t):=\overline{W}(t)+\varepsilon^{-1/2}\int_{0}^{t}P_{m(\varepsilon)}g(s)\,\mathrm{d}s$ is a cylindrical Wiener process on $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ and $\bar{\theta}_{\varepsilon,m(\varepsilon)}$ is a solution of
$\mathrm{d}\bar{\theta}_{\varepsilon,m(\varepsilon)}(t)=$ $-\Lambda^{2\alpha}\bar{\theta}_{\varepsilon,m(\varepsilon)}(t)\,\mathrm{d}t-P_{m}\left(u_{\bar{\theta}_{\varepsilon,m(\varepsilon)}}(t)\cdot\nabla\bar{\theta}_{\varepsilon,m(\varepsilon)}(t)\right)\,\mathrm{d}t$ (6.4)
$+\sqrt{\varepsilon}\Lambda^{2\beta}\,\mathrm{d}W^{m(\varepsilon)}(t)+P_{m(\varepsilon)}\Lambda^{2\beta}g\,\mathrm{d}t$
with the initial data $\bar{\theta}_{\varepsilon,m(\varepsilon)}(0)\sim\mathcal{G}(P_{m(\varepsilon)}\theta(0),\varepsilon P_{m(\varepsilon)}/2).$
Let $\pi_{\varepsilon,m(\varepsilon),\theta}$ be the law of $\bar{\theta}_{\varepsilon,m(\varepsilon)}$ on $\left(\Omega,\mathcal{F},\mathbb{P}\right)$. According to (6.3), the definition of relative entropy, and the expression of $\mathcal{I}_{0}$, it holds that
$\varepsilon\operatorname{Ent}\left(\pi_{\varepsilon,m(\varepsilon),\theta}|\mu_{\varepsilon,m(\varepsilon)}\right)$ $\leqslant\varepsilon\operatorname{Ent}(\mathbb{P}|\overline{\mathbb{P}})=\varepsilon\mathbb{E}\left[\log Z^{\varepsilon,m(\varepsilon)}(T)\right]$
$=\frac{1}{2}\int_{0}^{T}\|P_{m(\varepsilon)}g(s)\|_{H}^{2}\,\mathrm{d}s+\varepsilon\mathbb{E}\left[\log Y_{0}^{\varepsilon,m(\varepsilon)}\right]$
@@
For every $\varepsilon>0$ and $m(\varepsilon)\in\mathbb{N}_{+}$ satisfying the scaling regime (2.5), let $\theta,\,g,\,\pi^{\varepsilon,m(\varepsilon),\theta}$ and $\mu_{\varepsilon,m(\varepsilon)}$ be as in Lemma 6.2. Then $\pi^{\varepsilon,m(\varepsilon),\theta}$ converges to $\delta_{\theta}$ in law on $\mathbb{X}_{\alpha,\beta}$ as $\varepsilon\rightarrow 0$.
@@
The proof will be divided into three steps. First, we show that $\{\pi^{\varepsilon,m(\varepsilon),\theta}\}$ is tight in $\mathbb{X}_{\alpha,\beta}$ and that any limit point $\bar{\theta}$ is a weak solution of (1.1) in the sense of Definition 4.1. Next, we prove the $H^{\alpha-2\beta}$-energy inequality for $\bar{\theta}$. Finally, we conclude that $\bar{\theta}=\theta$ in $\mathbb{X}_{\alpha,\beta}$ by using the weak-strong uniqueness.